OpenStudy (anonymous):

What is the y-intercept of the line represented by the following equation? 5x-3y=-2 A.(0,2/3) B.(2/3,0) C.(0,-2/5) D.(-2/5,0) help please!!

OpenStudy (anonymous):

5x-3y=-2 i.e. 5x+2=3y or 3y = 5x+2 \[y= \frac{5}{3}x+{2}{3}\] comparing it with y=mx+ c we find y intercept =\[c=\frac{2}{3}\] Thus option A.(0,2/3) is correct Another way to find y-intercept is plug x=0 in the equation so we find 5*0-3y=-2 i.e. -3y =-2 i.e. y= 2/3 Thus (x,y) = (0, 2/3) is the ans

OpenStudy (anonymous):

thank you!